OpenStudy (anonymous):

There are three glasses containing juice. Glass 1 contains 160 ml of juice, while glass 2 contains 50 ml. If one fourth of the juice from glass 1 is poured into glass 2 and one fifth of the juice from glass 3 is poured into glass 2, all the three glasses will end up having the same amount of juice. How much juice is there in glass 3 to begin with?

ganeshie8 (ganeshie8):

so there are 3 glasses wid juice

ganeshie8 (ganeshie8):

to start wid, glass1 = 160 ml glass2 = 50 ml say, glass3 = x ml

ganeshie8 (ganeshie8):

one fourth of the juice from glass 1 is poured into glass 2 after that, how much is left in glass1 ?

ganeshie8 (ganeshie8):

@Got.that.power13 did u solve it already, you have closed the question :o

OpenStudy (anonymous):

Yes :) thanks

ganeshie8 (ganeshie8):

okie good to hear :) wat did u get for glass3 ?

OpenStudy (anonymous):

the amount of juice in glass 2 = (1/4 x 160) + 50 + (z/5) = 40 + 50 + z/5. Remaining juice in glass 1 = 160 - 40 = 120 ml. Since all glasses contain the same amount of juice after pouring, 120 = 90 + z/5, or 30 = z/5, or z = 150 ml

ganeshie8 (ganeshie8):

Awesome ! i got 150 too

OpenStudy (anonymous):

:D